The height of a triangle is 3cm less then the Length of the base. The Area of the triangle is 35〖cm〗^2. What is the height and the base?
Answer by shree840(260)

(x-3)x/2=35
x^2-3x=70
X^2-3x-70=0
(x-10)(x+7)=0
x=10 or -7
ignoring -7 (length can not be negative
x=10 height 7 base 10